OpenStudy (anonymous):

What is the EXACT area of the circle that has a radius of 1/2 inch?

OpenStudy (anonymous):

a. \[\frac{ 1 }{ 4 }\pi inches^{2}\] b.\[\frac{ 1 }{ 2 }\pi inches ^{2}\] c.\[\pi inches ^{2}\] d.\[2\pi inches^2\]

OpenStudy (anonymous):

http://mathforum.org/library/drmath/view/55433.html

OpenStudy (anonymous):

\[\large A=\pi \times r^{2}\]\[\large A=\pi \times (\frac{1}{2} \times \frac{1}{2})\]\[\large A=\pi \times \frac{1}{4}\]\[\large A=\frac{1}{4}\pi\]

OpenStudy (anonymous):

@lengendarysadist thank you so much

OpenStudy (anonymous):

\[\huge \color{aqua}N\color{fuchsia}o \space \color{lime}P \color{orange}r \color{blue}o \color{maroon}b \color{red}l \color{olive}e \color{purple}m \ddot\smile \]